modified output of getStatus command to show more information about selected AGW unit. It now shows unit number, probe position, associated rotator and filter name which closes IT #2473.
added basic functions and config parameter to AGw configuration files to support dark files in preparation to close IT #2531, this is not fully implemented yet but you might see some hints that it is coming soon
removed WFS_pupil_direction from config files and replaced it with WFS_pupil_dir_x and WFS_pupil_dir_y, to control the pupil shift direction on the axes independently. This was already in use as part of a patch to GCS
IT #2621 should be fixed in this release, which caused postage stamp readouts in ACQUIRE mode or for acquisition images with AGW #2 (it is actually a work-around for a problem in AzCam 4.0)
fixed a problem with Azcam 4.0 images not getting the FITS headers extension they were supposed to get. This problem also only applied to AGW #2 images.
extended the delay before use of AGw unit after power up from 6 to 10 seconds as I still experienced problems on the mountain when using the auto powerUp feature
side parameter is now case insensitive
Support Astronomer Overview Information
The PSF now puts ACTOPT messages in SysLog for active optics operations on the secondary mirror.
New IRC commands ClearHotspot and ClearOffset exist to remove the PresetTelescope hotspot and offset objects. Note that the hotspot and offsets objects are not yet supported by the TCS.
Specific Tests that need to be performed with the Telescope
The proper unit conversion was implemented for the Aftercooler pressure differential, HX0410. The new device reports in Inches of Water versus PSI, and the unit conversion to kPa was not updated at the time of the device change.
Operating System, System Configuration, TCS and Hardware Computers
Updated telemetry classes to support telemetry 9 and greater.
Added code to handle exception which can be thrown by accessing properties of the mirror while it is in a disconnected state.
Added an enable-disable to the communication for the adsc and terc. In the case of the secondary, the TCP connection is broken, but for the tertiary, commands are just rejected, as severing the connection for one side will affect the other, since both sides use one UMAC.
This can be set via an lbt.conf setting so that at startup the communication is disabled to the UMAC, or via a button on each of the GUIs during runtime.
A new bit has been added to reflective memory called 'is_comm_disabled' showing the state of the communication to the UMAC for both objects.
Fixed bug in file upload to PMAC which would allow the file upload to continue, even if there was an error.
Added code to allow the OSS GUI to sense if the build it belongs to is not the same as the current build. The windows will be turned yellow if this is the case.
Added 'reverse_sel' to the terc selector to allow for the fact its direction may be reversed. This setting is in OSS.conf.
Removed the 'Version.hpp' file, since the version information is now held in 'Main.cpp'.
Paolo from ADS has modified the cfg file for the tertiary. the original name is 'UMAC_config_file_right_arm_fixed.CFG'. He has made changes to these files based on calibrations done to the UMAC on the mountain, since the amplifiers are new.
Under Paolo's direction, the focal station positions have been modified for the right side. The sign may need to be reversed in OSS.conf to be consistent with what is expected for the selector motion.
Under Paolo's instruction, modified the indicators for the 'isMirrorHomed' and 'isSelectorHomed'. Also fixed homing offsets for the home sequences for the selector and mirror.
Added more descriptive text to the state and status transistion events, as well as fixed the color on the Main OSS GUI.
Modified the 'wait for in position' code to fix a bug which caused it to return too fast.
Added more verbose error messages to the UMAC. These will be bubbled up to the IIF if an error occurs while running a command.
Finish() is no longer called on the telemetry collectors so the subsystem will not hang at exit if telemetry cannot flush the buffers. Any uncommitted samples in the buffers will be discarded.
The primary mirror position limits are tightened close to the horizon (> 0.3 degree elevation) to avoid some panics.
The secondary and tertiary mirrors now have position limits similar to the primary mirror. Variables in lbt.conf specify the various limit values, including whether the checks are active. The checks are currently disabled.
SysLog messages were added for active optics operations on the secondary mirror.